# Mirror and Preconditioned Gradient Descent in Wasserstein Space

### v1

@@ 1 Introduction
Minimizing functionals on the space of probability distributions has become ubiquitous in Machine Learning for e.g. sampling [12, 118], generative modeling [13] or learning neural networks [33, 83], and is a challenging task as it is an infinite-dimensional problem. Wasserstein gradient flows [4] provide an elegant way to solve such problems on the Wasserstein space, i.e., the space of probability distributions with bounded second moment, equipped with the 2-Wasserstein distance from optimal transport (OT). These flows provide continuous paths of distributions decreasing the objective functional and can be seen as analog to Euclidean gradient flows [103]. Their implicit time discretization, referred to as the JKO scheme [61], has been studied in depth [88, 1, 25, 103]. In contrast, explicit schemes, despite being easier to implement, have been less investigated. Most previous works focus on the optimization of a specific objective functional with a time-discretation of its gradient flow with the 2-Wasserstein metrics. For instance, the forward Euler discretization leads to the Wasserstein gradient descent. The latter takes the form of gradient descent (GD) on the position of particles for functionals with a closed-form over discrete measures, e.g. Maximum Mean Discrepancy (MMD), which can be of interest to train neural networks [6]. For objectives involving absolutely continuous measures, such as the Kullback-Leibler (KL) divergence for sampling, other discretizations can be easily computed such as the Unadjusted Langevin Algorithm (ULA) [99]. This leaves the question open of assessing the theoretical and empirical performance of other optimization algorithms relying on alternative geometries and time-discretizations.
@@
Mirror descent has already attracted attention in the sampling community, and some popular algorithms have been extended in this direction. For instance, ULA was adapted into the Mirror Langevin algorithm [58, 120, 30, 59, 3, 72]. Other sampling algorithms have received their counterpart mirror versions such as the Metropolis Adjusted Langevin Algorithm [108], diffusion models [75], Stein Variational Gradient Descent (SVGD) [106], or even Wasserstein gradient descent [105]. Preconditioned Wasserstein gradient descent has been also recently proposed for specific geometries in [41, 29] to minimize the KL in a more efficient way, but without an analysis in discrete time. All the previous references focus on optimizing the KL as an objective, while Wasserstein gradient flows have been studied in machine learning for different functionals such as more general $f$-divergences [5, 86], interaction energies [71], MMDs [6, 67, 56, 55] or Sliced-Wasserstein (SW) distances [18, 79, 42, 15]. In this work, we propose to bridge this gap by providing a general convergence theory of both mirror and preconditioned gradient descent schemes for general target functionals, and investigate as well empirical benefits of alternative transport geometries for optimizing functionals on the Wasserstein space. We emphasize that the latter is different from [8, 62], wherein mirror descent is defined in the Radon space of probability distributions, using the flat geometry defined by TV or $L^{2}$ norms on measures, see Appendix A for more details.
@@
We are interested in minimizing a functional $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}\cup\{+\infty\}$ over probability distributions, through schemes of the form, for $k\geq 0$,
@@
with different costs $\mathrm{d}:L^{2}(\mu_{k})\times L^{2}(\mu_{k})\to\mathbb{R}_{+}$, and in providing convergence conditions. While we can recover a map $\bar{\mathrm{T}}=\mathrm{T}_{k}\circ\mathrm{T}_{k-1}\dots\circ\mathrm{T}_{1}$ such that $\mu_{k}=\bar{\mathrm{T}}_{\#}\mu_{0}$, the scheme (1) proceeds by successive regularized linearizations retaining the Wasserstein structure, since the tangent space to $\mathcal{P}_{2}(\mathbb{R}^{d})$ at $\mu$ is a subset of $L^{2}(\mu)$ [89]. This paper is organized as follows. In Section 2, we provide some background on Bregman divergences and differentiability over the Wasserstein space. In Section 3, we consider Bregman divergences on $L^{2}(\mu)$ for the cost in (1), generalizing the mirror descent scheme to the Wasserstein space. In Section 4, we consider alternative costs in (1), that are analogous to OT distances with translation-invariant cost, extending the dual space preconditioning scheme to the latter space. Finally, in Section 5, we apply the two schemes to different objective functionals, including standard free energy functionals such as interaction energies and KL divergence, but also to Sinkhorn divergences [47] or SW [95, 16] with polynomial preconditioners on single-cell datasets.
@@
Notation. Consider the set $\mathcal{P}_{2}(\mathbb{R}^{d})$ of probability measures $\mu$ on $\mathbb{R}^{d}$ with finite second moment and $\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})\subset\mathcal{P}_{2}(\mathbb{R}^{d})$ its subset of absolutely continuous probability measures with respect to the Lebesgue measure. For any $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, we denote by $L^{2}(\mu)$ the Hilbert space of functions $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $\int\|f\|^{2}\mathrm{d}\mu<\infty$ equipped with the norm $\|\cdot\|_{L^{2}(\mu)}$ and inner product $\langle\cdot,\cdot\rangle_{L^{2}(\mu)}$. For a Hilbert space $X$, the Fenchel transform of $f:X\to\mathbb{R}$ is $f^{*}(y)=\sup_{x\in X}\ \langle x,y\rangle-f(x)$. Given a measurable map $\mathrm{T}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $\mathrm{T}_{\#}\mu$ is the pushforward measure of $\mu$ by $\mathrm{T}$; and $\mathrm{T}\star\mu=\int\mathrm{T}(\cdot-x)\mathrm{d}\mu(x)$. For $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the 2-Wasserstein distance is $\mathrm{W}_{2}^{2}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\int\|x-y\|^{2}\ \mathrm{d}\gamma(x,y)$, where $\Pi(\mu,\nu)$ is the set of couplings between $\mu$ and $\nu$, and we denote by $\Pi_{o}(\mu,\nu)$ the set of optimal couplings. We refer to the metric space $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathrm{W}_{2})$ as the Wasserstein space.
@@ 2 Background
Frigyik et al. [50, Definition 2.1] defined the Bregman divergence of Fréchet differentiable functionals. In our case, we only need Gâteaux differentiability. In this paper, $\nabla$ refers to the Gâteaux differential, which coincides with the Fréchet derivative if the latter exists.
@@
Let $\psi_{\mu},\phi_{\mu}:L^{2}(\mu)\to\mathbb{R}$ convex and continuously Gâteaux differentiable. We say that $\psi_{\mu}$ is $\beta$-smooth (respectively $\alpha$-convex) relative to $\phi$ if and only if for all $\mathrm{T},\mathrm{S}\in L^{2}(\mu),\ \mathrm{d}_{\psi_{\mu}}(\mathrm{T},\mathrm{S})\leq\beta\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{S})$ (respectively $\mathrm{d}_{\psi_{\mu}}(\mathrm{T},\mathrm{S})\geq\alpha\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{S})$).
Similarly to the Euclidean case [81], relative smoothness and convexity are equivalent with respectively $\beta\phi_{\mu}-\psi_{\mu}$ and $\psi_{\mu}-\alpha\phi_{\mu}$ being convex (see Section B.2). Yet, proving the convergence of (1) requires only that these properties hold at specific functions (directions), a fact we will soon exploit.
In some situations, we need the $L^{2}$ Fenchel transform $\phi_{\mu}^{*}$ of $\phi_{\mu}$ to be differentiable, e.g. to compute its Bregman divergence $\mathrm{d}_{\phi_{\mu}^{*}}$. We show in Lemma 16 that a sufficient condition to satisfy this property is for $\phi_{\mu}$ to be strictly convex, lower semicontinuous and superlinear, i.e. $\lim_{\|\mathrm{T}\|\to\infty}\phi_{\mu}(\mathrm{T})/\|\mathrm{T}\|_{L^{2}(\mu)}=+\infty$. Moreover, in this case, $(\nabla\phi_{\mu})^{-1}=\nabla\phi_{\mu}^{*}$. When needed, we will suppose that $\phi_{\mu}$ satisfies this.
@@
The definitions of Bregman divergences on $L^{2}(\mu)$ and of $\mathrm{W}_{2}$-differentiability enable us to consider alternative Bregman potentials than the $L^{2}(\mu)$-norm mentioned above. For instance, for $V$ and $W$ convex, differentiable and $L$-smooth with $W$ even, we can use potential energies $\phi_{\mu}^{V}(\mathrm{T}):=\mathcal{V}(\mathrm{T}_{\#}\mu)$, for which $\mathrm{d}_{\phi_{\mu}^{V}}(\mathrm{T},\mathrm{S})=\int\mathrm{d}_{V}\big{(}\mathrm{T}(x),\mathrm{S}(x)\big{)}\mathrm{d}\mu(x)$ where $\mathrm{d}_{V}$ is the Bregman divergence of $V$ on $\mathbb{R}^{d}$. Notice that $\phi_{\mu}(\mathrm{T})=\frac{1}{2}\|\mathrm{T}\|^{2}_{L^{2}(\mu)}$ is a specific example of a potential energy where $V=\frac{1}{2}\|\cdot\|^{2}$. In particular, we have $\nabla_{\mathrm{W}_{2}}\mathcal{V}(\mu)=\nabla V$. We will also consider interaction energies $\phi_{\mu}^{W}(\mathrm{T}):=\mathcal{W}(\mathrm{T}_{\#}\mu)$, for which $\mathrm{d}_{\phi_{\mu}^{W}}(\mathrm{T},\mathrm{S})=\iint\mathrm{d}_{W}\big{(}\mathrm{T}(x)-\mathrm{T}(x^{\prime}),\mathrm{S}(x)-\mathrm{S}(x^{\prime})\big{)}\mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$ (see Section I.3). In that case, $\nabla_{\mathrm{W}_{2}}\mathcal{W}(\mu)=\nabla W\star\mu$. We will also use $\phi_{\mu}^{\mathcal{H}}(\mathrm{T})=\mathcal{H}(\mathrm{T}_{\#}\mu)$ with $\mathcal{H}$ the negative entropy. Note that Bregman divergences on the Wasserstein space using these functionals were proposed by Li [73], but only for $\mathrm{S}=\mathrm{Id}$ and OT maps $\mathrm{T}$.
@@
For instance, free energies such as potential or interaction energies with convex $V$ or $W$, or the negative entropy, are convex along geodesics [102, Section 7.3]. However, some popular functionals, such as the 2-Wasserstein distance $\mu\mapsto\frac{1}{2}\mathrm{W}_{2}^{2}(\mu,\eta)$ itself, for a given $\eta\in\mathcal{P}_{2}(\mathbb{R}^{d})$, are not convex along geodesics. Instead Ambrosio et al. [4, Theorem 4.0.4] showed that it was sufficient for the convergence of the gradient flow to be convex along other curves, e.g. along particular generalized geodesics for the 2-Wasserstein distance [4, Lemma 9.2.7], which, for $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, are of the form $\mu_{t}=\big{(}(1-t)\mathrm{T}_{\eta}^{\mu}+t\mathrm{T}_{\eta}^{\nu}\big{)}_{\#}\eta$ for $\mathrm{T}_{\eta}^{\mu}$, $T_{\eta}^{\nu}$ OT maps from $\eta$ to $\mu$ and $\nu$. Observing that for $\phi_{\mu}(\mathrm{T})=\frac{1}{2}\|\mathrm{T}\|_{L^{2}(\mu)}^{2}$, we can rewrite (3) as $\alpha\mathrm{d}_{\phi_{\mu_{0}}}(\mathrm{T}_{\mu_{0}}^{\mu_{1}},\mathrm{Id})\leq\mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{0}}}(\mathrm{T}_{\mu_{0}}^{\mu_{1}},\mathrm{Id})$, we see that being convex along geodesics boils down to being convex in the $L^{2}$ sense for $\mathrm{S}=\mathrm{Id}$ and $\mathrm{T}$ chosen as an OT map. This observation motivates us to consider a more refined notion of convexity along curves.
@@
Notice that in contrast with Definition 2, Definition 3 is stated for a fixed distribution $\mu$ and directions ($\mathrm{S},\mathrm{T}$), and involves comparisons between Bregman divergences depending on $\mu$ and curves $(\mathrm{T}_{s})_{s\in[0,1]}$ depending on $\mathrm{S},\mathrm{T}$. The larger family of $\mathrm{S}$ and $\mathrm{T}$ for which Definition 3 holds, the more restricted is the notion of convexity of $\mathcal{F}-\alpha\mathcal{G}$ (resp. of $\beta\mathcal{G}-\mathcal{F}$) on $\mathcal{P}_{2}(\mathbb{R}^{d})$. For instance, 2-Wasserstein generalized geodesics with anchor $\eta\in\mathcal{P}_{2}(\mathbb{R}^{d})$ correspond to considering $\mathrm{S},\mathrm{T}$ as all the OT maps originating from $\eta$, among which geodesics are particular cases when taking $\eta=\mu$ (hence $\mathrm{S}=\mathrm{Id}$). If we furthermore ask for $\alpha$-convexity to hold for all $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\mathrm{T},\mathrm{S}\in L^{2}(\mu)$ (i.e., not only OT maps), then we recover the convexity along acceleration free-curves as introduced in [109, 91, 27]. Our motivation behind Definition 3 is that the convergence proofs of MD and preconditioned GD require relative smoothness and convexity properties to hold only along specific curves.
Mirror (MD) and preconditioned gradient descent (PGD) on $\mathbb{R}^{d}$.
These schemes read respectively as $\nabla\phi(x_{k+1})-\nabla\phi(x_{k})=-\tau\nabla f(x_{k})$ [11] and $y_{k+1}-y_{k}=-\tau\nabla h^{*}\big{(}\nabla g(y_{k})\big{)}$ [82], where the objectives $f,g$ and the regularizers $h,\phi$ are convex $C^{1}$ functions from $\mathbb{R}^{d}$ to $\mathbb{R}$. The algorithms are closely related since, using the Fenchel transform and setting $g=\phi^{*}$ and $h^{*}=f$, we see that, for $y=\nabla\phi(x)$, the two schemes are equivalent when permuting the roles of the objective and of the regularizer. For MD, convergence of $f$ is ensured if $f$ is both $\nicefrac{{1}}{{\tau}}$-smooth and $\alpha$-convex relative to $\phi$ [81, Theorem 3.1]. Concerning PGD, assuming that $h,g$ are Legendre, $g(y_{n})$ converges to the minimum of $g$ if $h^{*}$ is both $\nicefrac{{1}}{{\tau}}$-smooth and $\alpha$-convex relative to $g^{*}$ with $\alpha>0$ [82, Theorem 3.9].

### v2

@@ 1 Introduction
Minimizing functionals on the space of probability distributions has become ubiquitous in machine learning for e.g. sampling [13, 129], generative modeling [86, 53], learning neural networks [36, 90, 107], dataset transformation [4, 63], or modeling population dynamics [23, 120]. It is a challenging task as it is an infinite-dimensional problem. Wasserstein gradient flows [5] provide an elegant way to solve such problems on the Wasserstein space, i.e., the space of probability distributions with bounded second moment, equipped with the Wasserstein-2 distance from optimal transport (OT). These flows provide continuous paths of distributions decreasing the objective functional and can be seen as analog to Euclidean gradient flows [111]. Their implicit time discretization, referred to as the JKO scheme [66], has been studied in depth [95, 1, 26, 111]. In contrast, explicit schemes, despite being easier to implement, have been less investigated. Most previous works focus on the optimization of a specific objective functional with a time-discretation of its gradient flow with the Wasserstein-2 metrics. For instance, the forward Euler discretization leads to the Wasserstein gradient descent. The latter takes the form of gradient descent (GD) on the position of particles for functionals with a closed-form over discrete measures, e.g. Maximum Mean Discrepancy (MMD), which can be of interest to train neural networks [7, 30]. For objectives involving absolutely continuous measures, such as the Kullback-Leibler (KL) divergence for sampling, other discretizations can be easily computed such as the Unadjusted Langevin Algorithm (ULA) [106]. This leaves the question open of assessing the theoretical and empirical performance of other optimization algorithms relying on alternative geometries and time-discretizations.
@@
Mirror descent has already attracted attention in the sampling community, and some popular algorithms have been extended in this direction. For instance, ULA was adapted into the Mirror Langevin algorithm [62, 132, 32, 64, 3, 79, 121]. Other sampling algorithms have received their counterpart mirror versions such as the Metropolis Adjusted Langevin Algorithm [116], diffusion models [82], Stein Variational Gradient Descent (SVGD) [114], or even Wasserstein gradient descent [113]. Preconditioned Wasserstein gradient descent has been also recently proposed for specific geometries in [44, 31] to minimize the KL in a more efficient way, but without an analysis in discrete time. All the previous references focus on optimizing the KL as an objective, while Wasserstein gradient flows have been studied in machine learning for different functionals such as more general $f$-divergences [6, 93], interaction energies [78, 19], MMDs [7, 30, 72, 60, 59] or Sliced-Wasserstein (SW) distances [18, 86, 45, 15]. In this work, we propose to bridge this gap by providing a general convergence theory of both mirror and preconditioned gradient descent schemes for general target functionals, and investigate as well empirical benefits of alternative transport geometries for optimizing functionals on the Wasserstein space. We emphasize that the latter is different from [9, 67], wherein mirror descent is defined in the Radon space of probability distributions, using the flat geometry defined by TV or $L^{2}$ norms on measures, see Appendix A for more details.
@@
We are interested in minimizing a functional $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}\cup\{+\infty\}$ over probability distributions, through schemes of the form, for $\tau>0$, $k\geq 0$,
@@
with different costs $\mathrm{d}:L^{2}(\mu_{k})\times L^{2}(\mu_{k})\to\mathbb{R}_{+}$, and in providing convergence conditions. While we can recover a map $\bar{\mathrm{T}}=\mathrm{T}_{k}\circ\mathrm{T}_{k-1}\dots\circ\mathrm{T}_{1}$ such that $\mu_{k}=\bar{\mathrm{T}}_{\#}\mu_{0}$, the scheme (1) proceeds by successive regularized linearizations retaining the Wasserstein structure, since the tangent space to $\mathcal{P}_{2}(\mathbb{R}^{d})$ at $\mu$ is a subset of $L^{2}(\mu)$ [96]. This paper is organized as follows. In Section 2, we provide some background on Bregman divergences, as well as on differentiability and convexity over the Wasserstein space. In Section 3, we consider Bregman divergences on $L^{2}(\mu)$ for the cost in (1), generalizing the mirror descent scheme to the Wasserstein space. We study this new scheme by discussing its implementation, and proving its convergence under relative smoothness and convexity assumptions. In Section 4, we consider alternative costs in (1), that are analogous to OT distances with translation-invariant cost, extending the dual space preconditioning scheme to the latter space. Finally, in Section 5, we apply the two schemes to different objective functionals, including standard free energy functionals such as interaction energies and KL divergence, but also to Sinkhorn divergences [50] or SW [102, 16] with polynomial preconditioners on single-cell datasets.
@@
Notations. Consider the set $\mathcal{P}_{2}(\mathbb{R}^{d})$ of probability measures $\mu$ on $\mathbb{R}^{d}$ with finite second moment and $\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})\subset\mathcal{P}_{2}(\mathbb{R}^{d})$ its subset of absolutely continuous probability measures with respect to the Lebesgue measure. For any $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, we denote by $L^{2}(\mu)$ the Hilbert space of functions $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $\int\|f\|^{2}\mathrm{d}\mu<\infty$ equipped with the norm $\|\cdot\|_{L^{2}(\mu)}$ and inner product $\langle\cdot,\cdot\rangle_{L^{2}(\mu)}$. For a Hilbert space $X$, the Fenchel transform of $f:X\to\mathbb{R}$ is $f^{*}(y)=\sup_{x\in X}\ \langle x,y\rangle-f(x)$. Given a measurable map $\mathrm{T}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $\mathrm{T}_{\#}\mu$ is the pushforward measure of $\mu$ by $\mathrm{T}$; and $\mathrm{T}\star\mu=\int\mathrm{T}(\cdot-x)\mathrm{d}\mu(x)$. For $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the Wasserstein-2 distance is $\mathrm{W}_{2}^{2}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\int\|x-y\|^{2}\ \mathrm{d}\gamma(x,y)$, where $\Pi(\mu,\nu)=\{\gamma\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{d}),\ \pi^{1}_{\#}\gamma=\mu,\ \pi^{2}_{\#}\gamma=\nu\}$ with $\pi^{i}(x_{1},x_{2})=x_{i}$, is the set of couplings between $\mu$ and $\nu$, and we denote by $\Pi_{o}(\mu,\nu)$ the set of optimal couplings. When the optimal coupling is of the form $\gamma=(\mathrm{Id},\mathrm{T}_{\mu}^{\nu})_{\#}\mu$ with $\mathrm{Id}:x\mapsto x$ and $\mathrm{T}_{\mu}^{\nu}\in L^{2}(\mu)$ satisfying $(\mathrm{T}_{\mu}^{\nu})_{\#}\mu=\nu$, we call $\mathrm{T}_{\mu}^{\nu}$ the OT map. We refer to the metric space $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathrm{W}_{2})$ as the Wasserstein space. We note $S_{d}^{++}(\mathbb{R})$ the space of symmetric positive definite matrices, and for $x\in\mathbb{R}^{d}$, $\Sigma\in S_{d}^{++}(\mathbb{R})$, $\|x\|_{\Sigma}^{2}=x^{T}\Sigma x$.
@@ 2 Background
Frigyik et al. [54, Definition 2.1] defined the Bregman divergence of Fréchet differentiable functionals. In our case, we only need Gâteaux differentiability. In this paper, $\nabla$ refers to the Gâteaux differential, which coincides with the Fréchet derivative if the latter exists.
@@
Let $\psi_{\mu},\phi_{\mu}:L^{2}(\mu)\to\mathbb{R}$ be convex and continuously Gâteaux differentiable. We say that $\psi_{\mu}$ is $\beta$-smooth (respectively $\alpha$-convex) relative to $\phi_{\mu}$ if and only if for all $\mathrm{T},\mathrm{S}\in L^{2}(\mu),\ \mathrm{d}_{\psi_{\mu}}(\mathrm{T},\mathrm{S})\leq\beta\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{S})$ (respectively $\mathrm{d}_{\psi_{\mu}}(\mathrm{T},\mathrm{S})\geq\alpha\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{S})$).
Similarly to the Euclidean case [88], relative smoothness and convexity are equivalent to respectively $\beta\phi_{\mu}-\psi_{\mu}$ and $\psi_{\mu}-\alpha\phi_{\mu}$ being convex (see Section B.2). Yet, proving the convergence of (1) requires only that these properties hold at specific functions (directions), a fact we will soon exploit.
In some situations, we need the $L^{2}$ Fenchel transform $\phi_{\mu}^{*}$ of $\phi_{\mu}$ to be differentiable, e.g. to compute its Bregman divergence $\mathrm{d}_{\phi_{\mu}^{*}}$. We show in Lemma 18 that a sufficient condition to satisfy this property is for $\phi_{\mu}$ to be strictly convex, lower semicontinuous and superlinear, i.e. $\lim_{\|\mathrm{T}\|\to\infty}\phi_{\mu}(\mathrm{T})/\|\mathrm{T}\|_{L^{2}(\mu)}=+\infty$. Moreover, in this case, $(\nabla\phi_{\mu})^{-1}=\nabla\phi_{\mu}^{*}$. When needed, we will suppose that $\phi_{\mu}$ satisfies this assumption.
@@
Proposition 1.
The Wasserstein differentiable functionals include $c$-Wasserstein costs on $\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$ [74, Proposition 2.10 and 2.11], potential energies $\mathcal{V}(\mu)=\int V\mathrm{d}\mu$ or interaction energies $\mathcal{W}(\mu)=\frac{1}{2}\iint W(x-y)\ \mathrm{d}\mu(x)\mathrm{d}\mu(y)$ for $V:\mathbb{R}^{d}\to\mathbb{R}$ and $W:\mathbb{R}^{d}\to\mathbb{R}$ differentiable and with bounded Hessian [74, Section 2.4]. In particular, their Wasserstein gradients read as $\nabla_{\mathrm{W}_{2}}\mathcal{V}(\mu)=\nabla V$ and $\nabla_{\mathrm{W}_{2}}\mathcal{W}(\mu)=\nabla W\star\mu$. However, entropy functionals, e.g. the negative entropy defined as $\mathcal{H}(\mu)=\int\log\big(\rho(x)\big{missing})\mathrm{d}\mu(x)$ for distributions $\mu$ admitting a density $\rho$ w.r.t. the Lebesgue measure, are not Wasserstein differentiable. In this case, we can consider subgradients $\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu)$ at $\mu$ for which (2) becomes an inequality. To guarantee that the Wasserstein subgradient is not empty, we need $\rho$ to satisfy some Sobolev regularity, see e.g. [5, Theorem 10.4.13] or [108]. Then, if $\nabla\log\rho\in L^{2}(\mu)$, the only subgradient of $\mathcal{H}$ in the tangent space is $\nabla_{\mathrm{W}_{2}}\mathcal{H}(\mu)=\nabla\log\rho$, see [5, Theorem 10.4.17] and [47, Proposition 4.3]. Then, free energies are functionals that write as sums of potential, interaction and entropy terms [110, Chapter 7]. It is notably the case for the KL to a fixed target distribution, that is the sum of a potential and entropy term [129], or the MMD as a sum of a potential and interaction term [7].
@@
The definitions of Bregman divergences on $L^{2}(\mu)$ and of Wasserstein differentiability enable us to consider alternative Bregman potentials than the $L^{2}(\mu)$-norm mentioned above. For instance, for $V$ convex, differentiable and $L$-smooth, we can use potential energies $\phi_{\mu}^{V}(\mathrm{T}):=\mathcal{V}(\mathrm{T}_{\#}\mu)$, for which $\mathrm{d}_{\phi_{\mu}^{V}}(\mathrm{T},\mathrm{S})=\int\mathrm{d}_{V}\big{(}\mathrm{T}(x),\mathrm{S}(x)\big{)}\mathrm{d}\mu(x)$ where $\mathrm{d}_{V}$ is the Bregman divergence of $V$ on $\mathbb{R}^{d}$. Notice that $\phi_{\mu}(\mathrm{T})=\frac{1}{2}\|\mathrm{T}\|^{2}_{L^{2}(\mu)}$ is a specific example of a potential energy where $V=\frac{1}{2}\|\cdot\|^{2}$. Moreover, we will consider interaction energies $\phi_{\mu}^{W}(\mathrm{T}):=\mathcal{W}(\mathrm{T}_{\#}\mu)$ with $W$ convex, differentiable, $L$-smooth, and satisfying $W(-x)=W(x)$; for which $\mathrm{d}_{\phi_{\mu}^{W}}(\mathrm{T},\mathrm{S})=\frac{1}{2}\iint\mathrm{d}_{W}\big{(}\mathrm{T}(x)-\mathrm{T}(x^{\prime}),\mathrm{S}(x)-\mathrm{S}(x^{\prime})\big{)}\mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$ (see Section I.3). We will also use $\phi_{\mu}^{\mathcal{H}}(\mathrm{T})=\mathcal{H}(\mathrm{T}_{\#}\mu)$ with $\mathcal{H}$ the negative entropy. Note that Bregman divergences on the Wasserstein space using these functionals were proposed by Li [80], but only for $\mathrm{S}=\mathrm{Id}$ and optimal transport maps $\mathrm{T}$.
@@
For instance, free energies such as potential or interaction energies with convex $V$ or $W$, or the negative entropy, are convex along geodesics [110, Section 7.3]. However, some popular functionals, such as the Wasserstein-2 distance $\mu\mapsto\frac{1}{2}\mathrm{W}_{2}^{2}(\mu,\eta)$ itself, for a given $\eta\in\mathcal{P}_{2}(\mathbb{R}^{d})$, are not convex along geodesics. Instead Ambrosio et al. [5, Theorem 4.0.4] showed that it was sufficient for the convergence of the gradient flow to be convex along other curves, e.g. along particular generalized geodesics for the Wasserstein-2 distance [5, Lemma 9.2.7], which, for $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, are of the form $\mu_{t}=\big{(}(1-t)\mathrm{T}_{\eta}^{\mu}+t\mathrm{T}_{\eta}^{\nu}\big{)}_{\#}\eta$ for $\mathrm{T}_{\eta}^{\mu}$, $T_{\eta}^{\nu}$ OT maps from $\eta$ to $\mu$ and $\nu$. Observing that for $\phi_{\mu}(\mathrm{T})=\frac{1}{2}\|\mathrm{T}\|_{L^{2}(\mu)}^{2}$, we can rewrite (3) as $\alpha\mathrm{d}_{\phi_{\mu_{0}}}(\mathrm{T}_{\mu_{0}}^{\mu_{1}},\mathrm{Id})\leq\mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{0}}}(\mathrm{T}_{\mu_{0}}^{\mu_{1}},\mathrm{Id})$ and see that being convex along geodesics boils down to being convex in the $L^{2}$ sense for $\mathrm{S}=\mathrm{Id}$ and $\mathrm{T}$ chosen as an OT map. This observation motivates us to consider a more refined notion of convexity along curves.
@@
Notice that in contrast with Definition 2, Definition 3 is stated for a fixed distribution $\mu$ and directions ($\mathrm{S},\mathrm{T}$), and involves comparisons between Bregman divergences depending on $\mu$ and curves $(\mathrm{T}_{s})_{s\in[0,1]}$ depending on $\mathrm{S},\mathrm{T}$. The larger family of $\mathrm{S}$ and $\mathrm{T}$ for which Definition 3 holds, the more restricted is the notion of convexity of $\mathcal{F}-\alpha\mathcal{G}$ (resp. of $\beta\mathcal{G}-\mathcal{F}$) on $\mathcal{P}_{2}(\mathbb{R}^{d})$. For instance, Wasserstein-2 generalized geodesics with anchor $\eta\in\mathcal{P}_{2}(\mathbb{R}^{d})$ correspond to considering $\mathrm{S},\mathrm{T}$ as all the OT maps originating from $\eta$, among which geodesics are particular cases when taking $\eta=\mu$ (hence $\mathrm{S}=\mathrm{Id}$). If we furthermore ask for $\alpha$-convexity to hold for all $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\mathrm{T},\mathrm{S}\in L^{2}(\mu)$ (i.e., not only OT maps), then we recover the convexity along acceleration free-curves as introduced in [117, 98, 28]. Our motivation behind introducing Definition 3 is that the convergence proofs of MD and preconditioned GD require relative smoothness and convexity properties to hold only along specific curves.
Mirror descent and preconditioned gradient descent on $\mathbb{R}^{d}$.
These schemes read respectively as $\nabla\phi(x_{k+1})-\nabla\phi(x_{k})=-\tau\nabla f(x_{k})$ [12] and $y_{k+1}-y_{k}=-\tau\nabla h^{*}\big{(}\nabla g(y_{k})\big{)}$ [89], where the objectives $f,g$ and the regularizers $h,\phi$ are convex $C^{1}$ functions from $\mathbb{R}^{d}$ to $\mathbb{R}$. The algorithms are closely related since, using the Fenchel transform and setting $g=\phi^{*}$ and $h^{*}=f$, we see that, for $y=\nabla\phi(x)$, the two schemes are equivalent when permuting the roles of the objective and of the regularizer. For MD, convergence of $f$ is ensured if $f$ is both $\nicefrac{{1}}{{\tau}}$-smooth and $\alpha$-convex relative to $\phi$ [88, Theorem 3.1]. Concerning preconditioned GD, assuming that $h,g$ are Legendre, $\big{(}g(y_{k})\big{)}_{k}$ converges to the minimum of $g$ if $h^{*}$ is both $\nicefrac{{1}}{{\tau}}$-smooth and $\alpha$-convex relative to $g^{*}$ with $\alpha>0$ [89, Theorem 3.9].